# Topology of ultra-localized insulators and superconductors

### v1

@@ Acknowledgements.
The eigenvector matrix can be right-multiplied by a block-diagonal matrix $D$ containing $2\times 2$ blocks of the form $e^{i\phi_{\mathbf{n}}\tau_{3}}$ or $\tau_{1}e^{i\phi_{\mathbf{n}}\tau_{3}}$ without changing $H$. Physically, this freedom corresponds to the freedom to choose the phases of the electron-like wavefunctions. The phases of the hole-like wavefunctions are then determined by particle-hole symmetry. If the matrix $\tau_{1}$ appears in a block, the corresponding element $E_{\mathbf{n}}$ in the diagonal matrix $\Lambda$ must be multiplied by $-1$. Physically, this degree of freedom corresponds to an exchange of electron and hole labels, which must be accompanied by the change $E_{\mathbf{n}}\to-E_{\mathbf{n}}$. The relabeling matrix $P$ is proportional to $\tau_{0}$.
@@
As in class D, eigenvalues come in pairs ±$\pm E_{\mathbf{n}}$ and one may write
@@
where $U=\tau_{2}U^{*}\tau_{2}$ is a symplectic matrix and $\Lambda$ is a block-diagional matrix containing blocks $E_{\mathbf{n}}\tau_{3}$, where each eigenvalue pair ±$\pm E_{\mathbf{n}}$ appears once. The Hamiltonian $H$ is not changed if $U$ is right-multiplied by a block-diagonal matrix $D$ containing blocks of the form $\tau_{0}e^{i\phi_{\mathbf{n}}\tau_{3}}$ or $\tau_{2}e^{i\phi_{\mathbf{n}}\tau_{3}}$, where, in the latter case, $E_{\mathbf{n}}$ must be simultaneously multiplied by $-1$. The unitary matrix $U$ may also be right-multiplied by a local relabeling matrix $P$, which is proportional to $\tau_{0}$. Class C does not admit topological zero modes.
@@
The eigenvalues come in pairs ±$\pm E_{\mathbf{n}}$ and one may write
@@
where $\Lambda$ is a block-diagional matrix containing blocks $E_{\mathbf{n}}\tau_{1}$, such that each eigenvalue pair ±$\pm E_{\mathbf{n}}$ appears once. The matrix $U_{+}$ is a unitary matrix. The Hamiltonian $H$ remains unchanged if $U$ is right-multiplied by a block-diagonal “phase matrix” $D$ having $2\times 2$ diagonal blocks $D_{\mathbf{n}}$ equal to ±$\pm\tau_{0}$ or ±$\pm i\tau_{3}$, whereby $E_{\mathbf{n}}$ is simultaneously multiplied by $-1$ if $D_{\mathbf{n}}=i\tau_{3}$ or $-i\tau_{3}$. The unitary matrix $U$ may also be right-multiplied by a local relabeling matrix $P$, which is proportional to $\tau_{0}$. Class CI has no topological zero modes.
To construct the auxiliary Hamiltonian, one uses the matrix $U_{+}$ only. Since $U_{+}$ is a unitary matrix without further symmetries, the auxiliary Hamiltonian $H_{\mathrm{aux}}$ is in tenfold-way class AIII. We note that $H$ is unchanged if $U_{+}$ is right-multiplied by a diagonal matrix $D_{+}$ with diagonal elements $D_{+,\mathbf{n}}$ equal ±$\pm 1$ or ±$\pm i$, whereby one has to simultaneously multiply $E_{\mathbf{n}}$ by $-1$ if $D_{+,\mathbf{n}}=\pm i$. Since a relabeling matrix $P$ is proportional to $\tau_{0}$, one may define a relabeling matrix $P_{+}$ that acts on $U_{+}$ only. In $d=1$ one has $K_{\rm aux}^{\prime}(1)={\cal K}_{\rm AIII}(1)=\mathbb{Z}$, whereas $K_{\rm aux}^{\prime}(1)=0$ for $d>1$. It follows that
@@
Class DIII.— In class DIII, $H$ satisfies the chiral symmetry $H=-\tau_{3}H\tau_{3}$ and the particle-hole antisymmetry $H=-\tau_{2}\sigma_{2}H^{*}\tau_{2}\sigma_{2}$. The eigenvalues come in twofold degenerate (Kramers) pairs ±$\pm E_{\mathbf{n}}$ and one may write
@@
[…]
As in class CI, to construct the auxiliary Hamiltonian, one uses the matrix $U_{+}$ only. Since no symmetries are imposed on $U_{+}$, the auxiliary Hamiltonian $H_{\mathrm{aux}}$ is in tenfold-way class AIII. We note that $H$ is unchanged if $U_{+}$ is right-multiplied by a diagonal matrix $D_{+}$ with $2\times 2$ diagonal blocks $D_{+,\mathbf{n}}$ equal ±$\pm\sigma_{\mathbf{n}}$ or ±$\pm i\sigma_{\mathbf{n}}$, whereby one has to simultaneously multiply $E_{\mathbf{n}}$ by $-1$ if $D_{+,\mathbf{n}}$ contains a factor $i$. Since a relabeling matrix $P$ is proportional to $\tau_{0}\sigma_{0}$, one may define a relabeling matrix $P_{+}$ proportional to $\sigma_{0}$ that acts on $U_{+}$ only. Since $P_{+}$ is proprortional to $\sigma_{0}$, in $d=1$ one has $K_{\rm aux}^{\prime}(1)=2\mathbb{Z}$, whereas ${\cal K}_{\rm AIII}=\mathbb{Z}$, so that
@@
class $\,d=1$ $d=2$ $d=3$ $d=4$ $d=5$ $d=6$ $d=7$ $d=8$ $d=9$
A 00 ${\text{Ch}_{1}}$ 00 ${\text{Ch}_{2}}$ 00 ${\text{Ch}_{3}}$ 00 ${\text{Ch}_{4}}$ 00
AIII ${\text{Wi}_{1}}$ 00 ${\text{Wi}_{3}}$ 00 ${\text{Wi}_{5}}$ 00 ${\text{Wi}_{7}}$ 00 ${\text{Wi}_{9}}$
AI 00 00 00 $2{\text{Ch}_{2}}$ 00 $\beta_{\text{Ch}_{3}}$ $\alpha_{\text{Ch}_{3}}$ ${\text{Ch}_{3}}$ 00
BDI ${\text{Wi}_{1}}$ 00 00 00 $2{\text{Wi}_{5}}$ 00 $\beta_{\text{Wi}_{9}}$ $\alpha_{\text{Wi}_{9}}$ ${\text{Wi}_{9}}$
D $\alpha_{\text{Ch}_{1}}$ ${\text{Ch}_{1}}$ 00 00 00 $2{\text{Ch}_{3}}$ 00 $\beta_{\text{Ch}_{5}}$ $\alpha_{\text{Ch}_{5}}$
DIII $\beta_{\text{Wi}_{3}}$ $\alpha_{\text{Wi}_{3}}$ ${\text{Wi}_{3}}$ 00 00 00 $2{\text{Wi}_{7}}$ 00 $\beta_{\text{Wi}_{11}}$
AII 00 $\beta_{\text{Ch}_{2}}$ $\alpha_{\text{Ch}_{2}}$ ${\text{Ch}_{2}}$ 00 00 00 $2{\text{Ch}_{4}}$ 00
CII $2{\text{Wi}_{1}}$ 00 $\beta_{\text{Wi}_{5}}$ $\alpha_{\text{Wi}_{5}}$ ${\text{Wi}_{5}}$ 00 00 00 $2{\text{Wi}_{9}}$
C 00 $2{\text{Ch}_{1}}$ 00 $\beta_{\text{Ch}_{3}}$ $\alpha_{\text{Ch}_{3}}$ ${\text{Ch}_{3}}$ 00 00 00
CI 00 00 $2{\text{Wi}_{3}}$ 00 $\beta_{\text{Wi}_{7}}$ $\alpha_{\text{Wi}_{7}}$ ${\text{Wi}_{7}}$ 00 00
@@
The topological invariants of conventional topological insulators and superconductors are well known Ryu et al. (2010). They are summarized in Table 3. The topological invariants are of integer ($\mathbb{Z}$) or of even-odd ($\mathbb{Z}_{2}$) type. The integer invariant is the Chern number ${\text{Ch}_{n}}$ or a higher-dimensional generalization of it for even dimensions $d=2n$ or it is the winding number ${\text{Wi}_{2n-1}}$ or a higher-dimensional generalization for odd dimensions $d=2n-1$. For the complex classes A and AIII the topological invariant is always an integer. For the real classes, there is one series of integer invariants which can be even or odd and one series of integer invariants that can take even values only. In addition, there are two series of $\mathbb{Z}_{2}$-type descendant invariants, which derive from the former series of integer invariants. The first (second) descendant $\alpha_{X}$ ($\beta_{X}$) is equal to the parity of the parent integer invariant $X$ in the same tenfold-way class, but in one (two) spatial dimension(s) higher, see Tab. 3.
@@
AIII ${\text{Wi}_{1}}$ 00 $({\text{Wi}_{3}},{\text{Wi}_{3}}^{\prime})$ 00 $({\text{Wi}_{5}},{\text{Wi}_{5}}^{\prime})$ 00 $({\text{Wi}_{7}},{\text{Wi}_{7}}^{\prime})$ 00 $({\text{Wi}_{9}},{\text{Wi}_{9}}^{\prime})$
A 00 00 ${\text{Wi}_{3}}$ 00 ${\text{Wi}_{5}}$ 00 ${\text{Wi}_{7}}$ 00 ${\text{Wi}_{9}}$
DIII $\beta_{\text{Wi}_{3}}$ 00 ${\text{Wi}_{3}}$ 00 ${\text{Wi}_{5}}$ 00 ${\text{Wi}_{7}}$ 00 ${\text{Wi}_{9}}$
CI 00 00 ${\text{Wi}_{3}}$ 00 ${\text{Wi}_{5}}$ 00 ${\text{Wi}_{7}}$ 00 ${\text{Wi}_{9}}$
BDI ${\text{Wi}_{1}}$ 00 00 00 $(2{\text{Wi}_{5}},2{\text{Wi}_{5}}^{\prime})$ 00 $(\beta_{\text{Wi}_{9}},\beta_{\text{Wi}_{9}}^{\prime})$ $(\alpha_{\text{Wi}_{9}},\alpha_{\text{Wi}_{9}}^{\prime})$ $({\text{Wi}_{9}},{\text{Wi}_{9}}^{\prime})$
AI 00 00 00 00 $2{\text{Wi}_{5}}$ 00 $\beta_{\text{Wi}_{9}}$ $\alpha_{\text{Wi}_{9}}$ ${\text{Wi}_{9}}$
D $\alpha_{\text{Ch}_{1}}$ 00 00 00 $2{\text{Wi}_{5}}$ 00 $\beta_{\text{Wi}_{9}}$ $\alpha_{\text{Wi}_{9}}$ ${\text{Wi}_{9}}$
CII $2{\text{Wi}_{1}}$ 00 $(\beta_{\text{Wi}_{5}},\beta_{\text{Wi}_{5}}^{\prime})$ $(\alpha_{\text{Wi}_{5}},\alpha_{\text{Wi}_{5}}^{\prime})$ $({\text{Wi}_{5}},{\text{Wi}_{5}}^{\prime})$ 00 00 00 $(2{\text{Wi}_{9}},2{\text{Wi}_{9}}^{\prime})$
AII 00 00 $\beta_{\text{Wi}_{5}}$ $\alpha_{\text{Wi}_{5}}$ ${\text{Wi}_{5}}$ 00 00 00 $2{\text{Wi}_{9}}$
C 00 00 $\beta_{\text{Wi}_{5}}$ $\alpha_{\text{Wi}_{5}}$ ${\text{Wi}_{5}}$ 00 00 00 $2{\text{Wi}_{9}}$
@@
class $\,d=1$ $d=2$ $d=3$ $d=4$ $d=5$ $d=6$ $d=7$ $d=8$ $d=9$
A 00 00 00 00 00 00 00 00 00
AIII ${\text{Wi}_{1}}$ 00 ${\text{Wi}_{3}}^{\prime}-{\text{Wi}_{3}}$ 00 ${\text{Wi}_{5}}^{\prime}-{\text{Wi}_{5}}$ 00 ${\text{Wi}_{7}}^{\prime}-{\text{Wi}_{7}}$ 00 ${\text{Wi}_{9}}^{\prime}-{\text{Wi}_{9}}$
AI 00 00 00 00 00 00 00 00 00
BDI ${\text{Wi}_{1}}$ 00 00 00 $2{\text{Wi}_{5}}^{\prime}-2{\text{Wi}_{5}}$ 00 $\beta_{\text{Wi}_{9}}^{\prime}+\beta_{\text{Wi}_{9}}$ $\alpha_{\text{Wi}_{9}}^{\prime}-\alpha_{\text{Wi}_{9}}$ ${\text{Wi}_{9}}^{\prime}-{\text{Wi}_{9}}$
D $\alpha_{\text{Ch}_{1}}$ 00 00 00 00 00 00 $\alpha_{\text{Wi}_{9}}$ ${\text{Wi}_{9}}\mod 2$
DIII $\beta_{\text{Wi}_{3}}$ 00 $2{\text{Wi}_{3}}$ 00 00 00 $2{\text{Wi}_{7}}$ 00 ${\text{Wi}_{9}}\mod 2$
AII 00 00 00 00 00 00 00 00 00
CII $2{\text{Wi}_{1}}$ 00 $\beta_{\text{Wi}_{5}}^{\prime}-\beta_{\text{Wi}_{5}}$ $\alpha_{\text{Wi}_{5}}^{\prime}-\alpha_{\text{Wi}_{5}}$ ${\text{Wi}_{5}}^{\prime}-{\text{Wi}_{5}}$ 00 00 00 $2{\text{Wi}_{9}}^{\prime}-2{\text{Wi}_{9}}$
C 00 00 00 $\alpha_{\text{Wi}_{5}}$ ${\text{Wi}_{5}}\mod 2$ 00 00 00 00
CI 00 00 $2{\text{Wi}_{3}}$ 00 ${\text{Wi}_{5}}\mod 2$ 00 $2{\text{Wi}_{7}}$ 00 00

### v2

@@ Acknowledgements.
The eigenvector matrix can be right-multiplied by a block-diagonal matrix $D$ containing $2\times 2$ blocks of the form $e^{i\phi_{\mathbf{n}}\tau_{3}}$ or $\tau_{1}e^{i\phi_{\mathbf{n}}\tau_{3}}$ without changing $H$. Physically, this freedom corresponds to the freedom to choose the phases of the electron-like wavefunctions. The phases of the hole-like wavefunctions are then determined by particle-hole symmetry. If the matrix $\tau_{1}$ appears in a block, the corresponding element $E_{\mathbf{n}}$ in the diagonal matrix $\Lambda$ must be multiplied by 1-1. Physically, this degree of freedom corresponds to an exchange of electron and hole labels, which must be accompanied by the change $E_{\mathbf{n}}\to-E_{\mathbf{n}}$. The relabeling matrix $P$ is proportional to $\tau_{0}$.
@@
As in class D, eigenvalues come in pairs $\pm E_{\mathbf{n}}$ and one may write
@@
where $U=\tau_{2}U^{*}\tau_{2}$ is a symplectic matrix and $\Lambda$ is a block-diagional matrix containing blocks $E_{\mathbf{n}}\tau_{3}$, where each eigenvalue pair $\pm E_{\mathbf{n}}$ appears once. The Hamiltonian $H$ is not changed if $U$ is right-multiplied by a block-diagonal matrix $D$ containing blocks of the form $\tau_{0}e^{i\phi_{\mathbf{n}}\tau_{3}}$ or $\tau_{2}e^{i\phi_{\mathbf{n}}\tau_{3}}$, where, in the latter case, $E_{\mathbf{n}}$ must be simultaneously multiplied by 1-1. The unitary matrix $U$ may also be right-multiplied by a local relabeling matrix $P$, which is proportional to $\tau_{0}$. Class C does not admit topological zero modes.
@@
The eigenvalues come in pairs $\pm E_{\mathbf{n}}$ and one may write
@@
where $\Lambda$ is a block-diagional matrix containing blocks $E_{\mathbf{n}}\tau_{1}$, such that each eigenvalue pair $\pm E_{\mathbf{n}}$ appears once. The matrix $U_{+}$ is a unitary matrix. The Hamiltonian $H$ remains unchanged if $U$ is right-multiplied by a block-diagonal “phase matrix” $D$ having $2\times 2$ diagonal blocks $D_{\mathbf{n}}$ equal to $\pm\tau_{0}$ or $\pm i\tau_{3}$, whereby $E_{\mathbf{n}}$ is simultaneously multiplied by 1-1 if $D_{\mathbf{n}}=i\tau_{3}$ or $i\tau_{3}$. The unitary matrix $U$ may also be right-multiplied by a local relabeling matrix $P$, which is proportional to $\tau_{0}$. Class CI has no topological zero modes.
To construct the auxiliary Hamiltonian, one uses the matrix $U_{+}$ only. Since $U_{+}$ is a unitary matrix without further symmetries, the auxiliary Hamiltonian $H_{\mathrm{aux}}$ is in tenfold-way class AIII. We note that $H$ is unchanged if $U_{+}$ is right-multiplied by a diagonal matrix $D_{+}$ with diagonal elements $D_{+,\mathbf{n}}$ equal $\pm 1$ or $\pm i$, whereby one has to simultaneously multiply $E_{\mathbf{n}}$ by 1-1 if $D_{+,\mathbf{n}}=\pm i$. Since a relabeling matrix $P$ is proportional to $\tau_{0}$, one may define a relabeling matrix $P_{+}$ that acts on $U_{+}$ only. In $d=1$ one has $K_{\rm aux}^{\prime}(1)={\cal K}_{\rm AIII}(1)=\mathbb{Z}$, whereas $K_{\rm aux}^{\prime}(1)=0$ for $d>1$. It follows that
@@
Class DIII.— In class DIII, $H$ satisfies the chiral symmetry $H=-\tau_{3}H\tau_{3}$ and the particle-hole antisymmetry $H=-\tau_{2}\sigma_{2}H^{*}\tau_{2}\sigma_{2}$. The eigenvalues come in twofold degenerate (Kramers) pairs $\pm E_{\mathbf{n}}$ and one may write
@@
Here $\Lambda$ is a block-diagonal matrix containing blocks $E_{\mathbf{n}}\tau_{1}\sigma_{0}$, such that each Kramers pair $\pm E_{\mathbf{n}}$ appears once. The matrix $U_{+}$ is a unitary matrix without further symmetries imposed. The Hamiltonian $H$ remains unchanged if $U$ is right-multiplied by a block-diagonal “phase matrix” $D$ having $4\times 4$ diagonal blocks $D_{\mathbf{n}}$ equal to $\pm\tau_{0}\sigma_{\mathbf{n}}$ or $\pm i\tau_{3}\sigma_{\mathbf{n}}$ with $\sigma_{\mathbf{n}}\in\mbox{SU(2)}$, whereby $E_{\mathbf{n}}$ is simultaneously multiplied by 1-1 if $D_{\mathbf{n}}$ contains $\tau_{3}$. The unitary matrix $U$ may also be right-multiplied by a local relabeling matrix $P$, which is proportional to $\tau_{0}\sigma_{0}$. Class DIII admits twofold (Kramers) degenerate topological zero modes. Hamiltonians $H$ with such topological zero-energy Kramers pairs must be excluded when defining topological ultra-localized superconductor in class DIII.
As in class CI, to construct the auxiliary Hamiltonian, one uses the matrix $U_{+}$ only. Since no symmetries are imposed on $U_{+}$, the auxiliary Hamiltonian $H_{\mathrm{aux}}$ is in tenfold-way class AIII. We note that $H$ is unchanged if $U_{+}$ is right-multiplied by a diagonal matrix $D_{+}$ with $2\times 2$ diagonal blocks $D_{+,\mathbf{n}}$ equal $\pm\sigma_{\mathbf{n}}$ or $\pm i\sigma_{\mathbf{n}}$, whereby one has to simultaneously multiply $E_{\mathbf{n}}$ by 1-1 if $D_{+,\mathbf{n}}$ contains a factor $i$. Since a relabeling matrix $P$ is proportional to $\tau_{0}\sigma_{0}$, one may define a relabeling matrix $P_{+}$ proportional to $\sigma_{0}$ that acts on $U_{+}$ only. Since $P_{+}$ is proprortional to $\sigma_{0}$, in $d=1$ one has $K_{\rm aux}^{\prime}(1)=2\mathbb{Z}$, whereas ${\cal K}_{\rm AIII}=\mathbb{Z}$, so that
@@
class d=1\,d=1 $d=2$ $d=3$ $d=4$ $d=5$ $d=6$ $d=7$ $d=8$ $d=9$
A 0 $\text{Ch}_{1}$} 0 $\text{Ch}_{2}$} 0 $\text{Ch}_{3}$} 0 $\text{Ch}_{4}$} 0
AIII $\text{Wi}_{1}$} 0 $\text{Wi}_{3}$} 0 $\text{Wi}_{5}$} 0 $\text{Wi}_{7}$} 0 $\text{Wi}_{9}$}
AI 0 0 0 $2{\text{Ch}_{2}}$ 0 $\beta_{\text{Ch}_{3}}$ $\alpha_{\text{Ch}_{3}}$ $\text{Ch}_{3}$} 0
BDI $\text{Wi}_{1}$} 0 0 0 $2{\text{Wi}_{5}}$ 0 $\beta_{\text{Wi}_{9}}$ $\alpha_{\text{Wi}_{9}}$ $\text{Wi}_{9}$}
D $\alpha_{\text{Ch}_{1}}$ $\text{Ch}_{1}$} 0 0 0 $2{\text{Ch}_{3}}$ 0 $\beta_{\text{Ch}_{5}}$ $\alpha_{\text{Ch}_{5}}$
DIII $\beta_{\text{Wi}_{3}}$ $\alpha_{\text{Wi}_{3}}$ $\text{Wi}_{3}$} 0 0 0 $2{\text{Wi}_{7}}$ 0 $\beta_{\text{Wi}_{11}}$
AII 0 $\beta_{\text{Ch}_{2}}$ $\alpha_{\text{Ch}_{2}}$ $\text{Ch}_{2}$} 0 0 0 $2{\text{Ch}_{4}}$ 0
CII $2{\text{Wi}_{1}}$ 0 $\beta_{\text{Wi}_{5}}$ $\alpha_{\text{Wi}_{5}}$ $\text{Wi}_{5}$} 0 0 0 $2{\text{Wi}_{9}}$
C 0 $2{\text{Ch}_{1}}$ 0 $\beta_{\text{Ch}_{3}}$ $\alpha_{\text{Ch}_{3}}$ $\text{Ch}_{3}$} 0 0 0
CI 0 0 $2{\text{Wi}_{3}}$ 0 $\beta_{\text{Wi}_{7}}$ $\alpha_{\text{Wi}_{7}}$ $\text{Wi}_{7}$} 0 0
@@
The topological invariants of conventional topological insulators and superconductors are well known Ryu et al. (2010). They are summarized in Table 3. The topological invariants are of integer ($\mathbb{Z}$) or of even-odd ($\mathbb{Z}_{2}$) type. The integer invariant is the Chern number $\text{Ch}_{n}$} or a higher-dimensional generalization of it for even dimensions $d=2n$ or it is the winding number $\text{Wi}_{2n-1}$} or a higher-dimensional generalization for odd dimensions d=2n1d=2n-1. For the complex classes A and AIII the topological invariant is always an integer. For the real classes, there is one series of integer invariants which can be even or odd and one series of integer invariants that can take even values only. In addition, there are two series of $\mathbb{Z}_{2}$-type descendant invariants, which derive from the former series of integer invariants. The first (second) descendant $\alpha_{X}$ ($\beta_{X}$) is equal to the parity of the parent integer invariant $X$ in the same tenfold-way class, but in one (two) spatial dimension(s) higher, see Tab. 3.
@@
AIII $\text{Wi}_{1}$} 0 $({\text{Wi}_{3}},{\text{Wi}_{3}}^{\prime})$ 0 $({\text{Wi}_{5}},{\text{Wi}_{5}}^{\prime})$ 0 $({\text{Wi}_{7}},{\text{Wi}_{7}}^{\prime})$ 0 $({\text{Wi}_{9}},{\text{Wi}_{9}}^{\prime})$
A 0 0 $\text{Wi}_{3}$} 0 $\text{Wi}_{5}$} 0 $\text{Wi}_{7}$} 0 $\text{Wi}_{9}$}
DIII $\beta_{\text{Wi}_{3}}$ 0 $\text{Wi}_{3}$} 0 $\text{Wi}_{5}$} 0 $\text{Wi}_{7}$} 0 $\text{Wi}_{9}$}
CI 0 0 $\text{Wi}_{3}$} 0 $\text{Wi}_{5}$} 0 $\text{Wi}_{7}$} 0 $\text{Wi}_{9}$}
BDI $\text{Wi}_{1}$} 0 0 0 $(2{\text{Wi}_{5}},2{\text{Wi}_{5}}^{\prime})$ 0 $(\beta_{\text{Wi}_{9}},\beta_{\text{Wi}_{9}}^{\prime})$ $(\alpha_{\text{Wi}_{9}},\alpha_{\text{Wi}_{9}}^{\prime})$ $({\text{Wi}_{9}},{\text{Wi}_{9}}^{\prime})$
AI 0 0 0 0 $2{\text{Wi}_{5}}$ 0 $\beta_{\text{Wi}_{9}}$ $\alpha_{\text{Wi}_{9}}$ $\text{Wi}_{9}$}
D $\alpha_{\text{Ch}_{1}}$ 0 0 0 $2{\text{Wi}_{5}}$ 0 $\beta_{\text{Wi}_{9}}$ $\alpha_{\text{Wi}_{9}}$ $\text{Wi}_{9}$}
CII $2{\text{Wi}_{1}}$ 0 $(\beta_{\text{Wi}_{5}},\beta_{\text{Wi}_{5}}^{\prime})$ $(\alpha_{\text{Wi}_{5}},\alpha_{\text{Wi}_{5}}^{\prime})$ $({\text{Wi}_{5}},{\text{Wi}_{5}}^{\prime})$ 0 0 0 $(2{\text{Wi}_{9}},2{\text{Wi}_{9}}^{\prime})$
AII 0 0 $\beta_{\text{Wi}_{5}}$ $\alpha_{\text{Wi}_{5}}$ $\text{Wi}_{5}$} 0 0 0 $2{\text{Wi}_{9}}$
C 0 0 $\beta_{\text{Wi}_{5}}$ $\alpha_{\text{Wi}_{5}}$ $\text{Wi}_{5}$} 0 0 0 $2{\text{Wi}_{9}}$
@@
class d=1\,d=1 $d=2$ $d=3$ $d=4$ $d=5$ $d=6$ $d=7$ $d=8$ $d=9$
A 0 0 0 0 0 0 0 0 0
AIII $\text{Wi}_{1}$} 0 ${\text{Wi}_{3}}^{\prime}-{\text{Wi}_{3}}$ 0 ${\text{Wi}_{5}}^{\prime}-{\text{Wi}_{5}}$ 0 ${\text{Wi}_{7}}^{\prime}-{\text{Wi}_{7}}$ 0 ${\text{Wi}_{9}}^{\prime}-{\text{Wi}_{9}}$
AI 0 0 0 0 0 0 0 0 0
BDI $\text{Wi}_{1}$} 0 0 0 $2{\text{Wi}_{5}}^{\prime}-2{\text{Wi}_{5}}$ 0 $\beta_{\text{Wi}_{9}}^{\prime}+\beta_{\text{Wi}_{9}}$ $\alpha_{\text{Wi}_{9}}^{\prime}-\alpha_{\text{Wi}_{9}}$ ${\text{Wi}_{9}}^{\prime}-{\text{Wi}_{9}}$
D $\alpha_{\text{Ch}_{1}}$ 0 0 0 0 0 0 $\alpha_{\text{Wi}_{9}}$ ${\text{Wi}_{9}}\mod 2$
DIII $\beta_{\text{Wi}_{3}}$ 0 $2{\text{Wi}_{3}}$ 0 0 0 $2{\text{Wi}_{7}}$ 0 ${\text{Wi}_{9}}\mod 2$
AII 0 0 0 0 0 0 0 0 0
CII $2{\text{Wi}_{1}}$ 0 $\beta_{\text{Wi}_{5}}^{\prime}-\beta_{\text{Wi}_{5}}$ $\alpha_{\text{Wi}_{5}}^{\prime}-\alpha_{\text{Wi}_{5}}$ ${\text{Wi}_{5}}^{\prime}-{\text{Wi}_{5}}$ 0 0 0 $2{\text{Wi}_{9}}^{\prime}-2{\text{Wi}_{9}}$
C 0 0 0 $\alpha_{\text{Wi}_{5}}$ ${\text{Wi}_{5}}\mod 2$ 0 0 0 0
CI 0 0 $2{\text{Wi}_{3}}$ 0 ${\text{Wi}_{5}}\mod 2$ 0 $2{\text{Wi}_{7}}$ 0 0